Qd=30-p Q's =2p


1
Expert's answer
2021-03-12T07:25:04-0500

The market equilibrium occurs when "Q_d=Q_s":


"30-p=2p,""30=3p,""p=\$10."

Finally, we can find the equilibrium quantity:


"Q_E=30-10=20."